# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ Appendix D Conjectures for ellipsoidal models
We conjecture that debiasing is possible in a more general settings than considered in this paper. Namely, one would like to consider the design matrix $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\bm{\Sigma}^{1/2}$ where $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ is non-singular, $\mathbf{Q}\in\mathbb{R}^{n\times n},\mathbf{O}\in\mathbb{R}^{p\times p}$ are orthogonal matrices and $\mathbf{D}\in\mathbb{R}^{n\times p}$ is diagonal matrix. We assume that $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ is observed and $\mathbf{O}$ is drawn uniformly from the orthogonal group $\mathbb{O}(p)$ independent of $\bm{\varepsilon},\mathbf{D},\mathbf{Q}$. We refer to this class of random design matrices as ellipsoidal invariant designs. The special case where $\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ is an isotropic Gaussian matrix is studied extensively in prior literature [12, 18, 54, 53, 11]. Similarly to the anisotropic Gaussian design case, the challenge in applying such a model arises when $\bm{\Sigma}$ is unknown. We discuss this in Remark D.2 at the end.
@@
$\hat{\bm{\beta}}^{u}=\hat{\bm{\beta}}+\frac{1}{\widehat{\mathsf{adj}}}\bm{\Sigma}^{-1}\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})$ (114)
@@
$\hat{\eta}_{*}(p):=\left(\frac{1}{p}\operatorname{Tr}\quantity(\widehat{\mathsf{adj}}\cdot\mathbf{I}_{p}+\bm{\Sigma}^{-1}\nabla^{2}\vec{h}(\hat{\bm{\beta}}))\right)^{-1}$ (116)
$\hat{\mathbf{{r}}}_{**}(p):=\hat{\bm{\beta}}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\bm{\Sigma}^{-1}\mathbf{X}^{\top}(\mathbf{X}\hat{\bm{\beta}}-\mathbf{y}),\quad\hat{\tau}_{**}(p):=\frac{\frac{1}{p}\left\|\mathbf{X}\hat{\mathbf{{r}}}_{**}-\mathbf{y}\right\|^{2}-\frac{n}{p}}{\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}}$
$\hat{\tau}_{*}(p):=\left(\frac{\hat{\eta}_{*}}{\widehat{\mathsf{adj}}}\right)^{2}\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{d_{i}^{2}}{\left(d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}}$
@@
The derivation of the above is by considering a change of variable $\tilde{\bm{\beta}}=\bm{\Sigma}^{1/2}\hat{\bm{\beta}}$ whereby $\tilde{\bm{\beta}}\in\underset{\mathbf{b}\in\mathbb{R}^{p}}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{b}\|^{2}+h\left(\bm{\Sigma}^{-\frac{1}{2}}\mathbf{b}\right)$ and using the iterates of the VAMP algorithm (for non-separable penalties [37, Algorithm 1]) to track $\tilde{\bm{\beta}}$. One can then obtain (116) and Conjecture D.1 from the state evolution of the VAMP algorithm [37, Eq. (19), Theorem 1]. If Conjecture D.1 holds, it will be straightforward to develop inference procedure for $\bm{\beta}^{\star}$. A main gap to prove Conjecture D.1 in our opinion is to establish an analogue of Proposition B.8, i.e. the non-separable VAMP iterates indeed tracks $\hat{\bm{\beta}}$. We leave the proof of Conjecture D.1 as an open problem.
Remark D.2.
@@
Appendix E Numerical experiments
E.1 Details of the design matrices
@@
Remark E.1 (Notations used in caption).
@@
Remark E.2 (Right-rotationally invariant).
@@
Remark E.3 (Comparison between designs in Figure 1 and Figure 3).
@@
Definition E.4 (Simulated Designs Specification).
@@
    $\mathsf{MatrixNormal}$-$\mathsf{B}$: $\mathbf{X}\sim N(\rm{0},\mathbf{\Sigma}^{\mathrm{(col)}}\otimes\mathbf{\Sigma}^{\mathrm{(row)}})$ where $\mathbf{\Sigma}^{\mathrm{(col)}}_{ij}=0.9^{|i-j|},\forall i,j\in[n]$ and $\bm{\Sigma}^{\mathrm{(row)}}\sim\mathsf{InverseWishart}\quantity(\mathbf{I}_{p},1.002\cdot p)$ (see Remark E.1 for notation);
@@
    $\mathsf{MultiCauchy}$: rows of $\mathbf{X}$ are sampled iid from $\mathsf{Mult}$-$\mathsf{t}(1,\mathbf{I}_{p})$ (see Remark E.1 for notation).
Definition E.5 (Real-data Designs Specification).
@@
E.2 QQ plots
Figure 5, 6 and 7 are QQ-plots of Figure 1, 3 top and bottom row experiments respectively.
@@ E.2 QQ plots
E.3 Marginal inference under exchangeability
Figure 8 below shows an illustration for this result focusing on =$\mathcal{I}=\{1\}$. Observe that we once again outperform degrees-of-freedom debiasing.
@@ E.3 Marginal inference under exchangeability
E.4 Alignment tests for simulated designs
@@
E.5 Misspecified setting
Our theory assumes that all relevant covariates are observed. In practice, there may be unobserved features $\mathbf{Z}$ and responses is generated from $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We found that our method is in fact relatively robust to such mis-specification. We demonstrate by introducing unobserved features $\mathbf{Z}$ of shape $n\times\frac{p}{2}$ (i.e. half as many as the observed features $\mathbf{X}\in\mathbb{R}^{n\times p}$) drawn from different distributions.
Figure 10: Z from MatrixNormal Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from MatrixNormal as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 11: Z from Spike. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from Spike as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 12: Z from LNN. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from LNN as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 13: Z from VAR. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from VAR as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 14: Z from Mult-t. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\epsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from Mult-t as described in Figure 1 with dimension half as many as $\mathbf{X}$.

### v6

@@ Appendix I Conjectures for Ellipsoidal Models
We conjecture that debiasing is possible in a more general settings than considered in this paper. Namely, one would like to consider the design matrix $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\bm{\Sigma}^{1/2}$ where $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ is non-singular, $\mathbf{Q}\in\mathbb{R}^{n\times n},\mathbf{O}\in\mathbb{R}^{p\times p}$ are orthogonal matrices and $\mathbf{D}\in\mathbb{R}^{n\times p}$ is diagonal matrix. We assume that $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ is observed and $\mathbf{O}$ is drawn uniformly from the orthogonal group $\mathbb{O}(p)$ independent of $\bm{\varepsilon},\mathbf{D},\mathbf{Q}$. We refer to this class of random design matrices as ellipsoidal invariant designs. The special case where $\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ is an isotropic Gaussian matrix is studied extensively in prior literature [13, 20, 62, 61, 12]. Similarly to the anisotropic Gaussian design case, the challenge in applying such a model arises when $\bm{\Sigma}$ is unknown. We discuss this in Remark I.2 at the end.
@@
[…]
@@
$\hat{\eta}_{*}(p)=\left(\frac{1}{p}\operatorname{Tr}\quantity(\widehat{\mathsf{adj}}\cdot\mathbf{I}_{p}+\bm{\Sigma}^{-1}\nabla^{2}\vec{h}(\hat{\bm{\beta}}))\right)^{-1}$ (175)
$\hat{\mathbf{{r}}}_{**}(p)=\hat{\bm{\beta}}+\frac{1}{\hat{\eta}_{*}-\widehat{\mathsf{adj}}}\bm{\Sigma}^{-1}\mathbf{X}^{\top}(\mathbf{X}\hat{\bm{\beta}}-\mathbf{y}),\quad\hat{\tau}_{**}(p)=\frac{\frac{1}{p}\left\|\mathbf{X}\hat{\mathbf{{r}}}_{**}-\mathbf{y}\right\|^{2}-\frac{n}{p}}{\frac{1}{p}\sum_{i=1}^{p}d_{i}^{2}}$
$\hat{\tau}_{*}(p)=\left(\frac{\hat{\eta}_{*}}{\widehat{\mathsf{adj}}}\right)^{2}\frac{1}{p}{{\sum}}_{i=1}^{p}\frac{d_{i}^{2}}{\left(d_{i}^{2}+\hat{\eta}_{*}-\widehat{\mathsf{adj}}\right)^{2}}$
@@
The derivation of the above is by considering a change of variable $\tilde{\bm{\beta}}=\bm{\Sigma}^{1/2}\hat{\bm{\beta}}$ whereby $\tilde{\bm{\beta}}\in\underset{\mathbf{b}\in\mathbb{R}^{p}}{\arg\min}\frac{1}{2}\|\mathbf{y}-\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{b}\|^{2}+h\left(\bm{\Sigma}^{-\frac{1}{2}}\mathbf{b}\right)$ and using the iterates of the VAMP algorithm (for non-separable penalties [44, Algorithm 1]) to track $\tilde{\bm{\beta}}$. One can then obtain (175) and Conjecture I.1 from the state evolution of the VAMP algorithm [44, Eq. (19), Theorem 1]. If Conjecture I.1 holds, it will be straightforward to develop inference procedure for $\bm{\beta}^{\star}$. A main gap to prove Conjecture I.1 in our opinion is to establish an analogue of Proposition C.10, i.e. the non-separable VAMP iterates indeed tracks $\hat{\bm{\beta}}$. We leave the proof of Conjecture I.1 as an open problem.
Remark I.2.
@@
Appendix J Numerical Experiments
J.1 Details of the design matrices
@@
Remark J.1 (Notations used in caption).
@@
Remark J.2 (Right-rotationally invariant).
@@
Remark J.3 (Comparison between designs in Figure 1 and Figure 3).
@@
Definition J.4 (Simulated Designs Specification).
@@
    $\mathsf{MatrixNormal}$-$\mathsf{B}$: $\mathbf{X}\sim N(\rm{0},\mathbf{\Sigma}^{\mathrm{(col)}}\otimes\mathbf{\Sigma}^{\mathrm{(row)}})$ where $\mathbf{\Sigma}^{\mathrm{(col)}}_{ij}=0.9^{|i-j|},\forall i,j\in[n]$ and $\bm{\Sigma}^{\mathrm{(row)}}\sim\mathsf{InverseWishart}\quantity(\mathbf{I}_{p},1.002\cdot p)$ (see Remark J.1 for notation);
@@
    $\mathsf{MultiCauchy}$: rows of $\mathbf{X}$ are sampled iid from $\mathsf{Mult}$-$\mathsf{t}(1,\mathbf{I}_{p})$ (see Remark J.1 for notation).
Definition J.5 (Real-data Designs Specification).
@@
J.2 QQ plots
Figure 5, 6 and 7 are $Q$-plots of Figure 1, 3 top and bottom row experiments respectively.
@@
J.3 Marginal inference under exchangeability
Figure 8 below shows an illustration for this result focusing on $\mathcal{I}=\{1\}$. Observe that we once again outperform degrees-of-freedom debiasing.
@@
J.4 Alignment tests for simulated designs
@@
J.5 Misspecified setting
Our theory assumes that all relevant covariates are observed. In practice, there may be unobserved features $\mathbf{Z}$ and responses is generated from $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\varepsilon}$. We found that our method is in fact relatively robust to such mis-specification. We demonstrate by introducing unobserved features $\mathbf{Z}$ of shape $n\times\frac{p}{2}$ (i.e. half as many as the observed features $\mathbf{X}\in\mathbb{R}^{n\times p}$) drawn from different distributions.
Figure 10: Z from MatrixNormal Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\varepsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from MatrixNormal as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 11: Z from Spike. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\varepsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from Spike as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 12: Z from LNN. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\varepsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from LNN as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 13: Z from VAR. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\varepsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from VAR as described in Figure 1 with dimension half as many as $\mathbf{X}$.
Figure 14: Z from Mult-t. Same settings and legends as Figure 3, bottom-left experiment, except that the response is generated from a misspecified model $\mathbf{y}=\mathbf{X}\bm{\beta}^{\star}+\mathbf{Z}\bm{\theta}^{\star}+\bm{\varepsilon}$. We set ${\theta}^{\star}_{i}\sim 0.2\cdot N(-5,1)+0.3\cdot N(2,1)+0.5\cdot\delta_{0}$. Here, we let $\mathbf{Z}$ be sampled from Mult-t as described in Figure 1 with dimension half as many as $\mathbf{X}$.